Please see the analysis of potential losses in its accounts receivable for Smith Industries below:

Age Past Due

Receivable Balance

Estimated Loss (%)

0-30 days past due

$   70,000

0.5%

31-60 days past due

45,000

2.5%

61-90 days past due

30,000

10.0%

Over three months past due

    5,000

95.0%

Total

$150,000

During the year, $1,500 of receivables were written off. The balance of Allowance for Uncollectible Accounts is a credit of $500 on December 31, 2021 prior to adjustments.

A.) Please calculate the bad debt expense for 2021.

B.) What is the amount of accounts receivable reported on Smith Industries' December 31, 2021 balance sheet?
Answer #1

Bad Debt expense = 70000 * 0.5% + 45000 * 2.5% + 30000 * 10% + 5000 * 95%

= 9225

Ending balance of allowance for doubtful accounts

= Balance before adjustments + Bad debt expense - Written off receivables

= 500 + 9225 - 1500

= 8225

Net account receivable shown in balance sheet = Gross receivables - Ending balance of allowance for doubtful account

= 150000 - 8225

= 141775